a boy pushes a 5000kg box along a floor with a force of 2000N.if the velocity of the box is uniform.What is the coefficient of friction between the box and the floor


1
Expert's answer
2022-03-13T18:46:12-0400

Since the speed is constant (zero acceleration) the friction force is equal to the pushing force by the second Newton's law:


"F_{fr} = 2000N"

On the other hand, the friction force is given as follows:


"F_{fr} = \mu N"

where "\mu" is the coefficient of friction, and "N" is the normal reation force, which is equal to the weight of the box, since the box does not move in vertical direction. Thus, obtain:


"F_{fr} = \mu mg"

where "m =5000kg" is the mass of the box, and "g = 9.8m\/s^2" is the gravitational acceleration. Finally, have:


"\mu = \dfrac{F_{fr}}{mg} = \dfrac{2000N}{5000kg\cdot 9.8m/s^2} \approx 0.04"

Answer. 0.04.